Explain the 3 different methods of accounting for Depreciation.  (Straight Line, Double Declining Balance, Units of Output or Production) Comment on the definition, calculation, journal entries, etc. of each method
Answer #1

Answer:

STRAIGHT- LINE DEPRECIATION (SLM)

This is the most easiest and generally acknowledged method of Depreciation. Also, it is most reasonable for resources that create income consistenly over their lives. what's more, the SLM technique give a fixed measure of deterioration for the entire time of the life of the advantage ignoring the utilization and effectiveness of the asset

.SLM annual depreciation charge = Depreciable Base / Estimated Service Life

Example; ASSET VALUE =1100000, SALVAGE VALUE = 720000, LIFE = 5 years,or 50000 Units

annual depreciation = 1100000-380000/5 = 144000

journal entry

Depreciation A/C Dr 144000

Accumulated Depreciation A/C Cr 144000

DECLINING BALANCE (DBM)

this is just an add on for Straight Line Depreciation method, and DBM calculates the depreciation disregarding the salvage value.

Here the double decline represents the depreciation charging twice than the Straight Line Depreciation method .

Same Example: 1ST year of depreciation % =1/5*2= 0.4 or 40%

=1100000*40% = 440000

journal entry

Depreciation A/C Dr 440000

Accumulated Depreciation A/C Cr 440000

UNIT OF PRODUCTION OR ACTIVITY METHOD

Unit of production is an asset by the amount that it is used. An assets can be measured by input units ( e.g., number of hours used) or output units (e.g., number of output produced).

Activity methord depreciation charge = depreciable base* units produced or hours used / total units of production or total hours usable over life

Same Example; it the asset actually produces the 9500 units in the first year

=720000*9500/50000 = 136800

journal entry

Depreciation A/C Dr 136800

Accumulated Depreciation A/C Cr 136800